Required information Problem 8-1A Plant asset costs; depreciation methods LO C1, P1 [The following information applies to the questions displayed below.] Timberly Construction makes a lump-sum purchase of several assets on January 1 at a total cash price of $840,000. The estimated market values of the purchased assets are building, $468,000; land, $302,250; land improvements, $29,250; and four vehicles, $175,500. Problem 8-1A Part 1-3 Required: 1-a. Allocate the lump-sum purchase price to the separate assets purchased. 1-b. Prepare the journal entry to record the purchase. 2. Compute the first-year depreciation expense on the building using the straight-line method, assuming a 15-year life and a $31,000 salvage value. 3. Compute the first-year depreciation expense on the land improvements assuming a five-year life and double-declining-balance depreciation.
Answer #1
1a
Allocation of total cost Appraised Value Percent of Total
Appraised Value
x Total cost of
Acquisition
Apportioned Cost
Building 468,000 48% x 840,000 403,200
Land 302,250 31% x 840,000 260,400
Land improvements 29,250 3% x 840,000 25,200
Vehicles 175,500 18% x 840,000 151,200
Total 975,000 100% 840,000
b
Date General Journal Debit Credit
1-Jan Building 403,200
Land 260,400
Land improvements 25,200
Vehicles 151,200
Cash 840,000
2
Depreciation expense on building 24813 =(403200-31000)/15
3
Depreciation rate 40% =1/5*2
Depreciation expense on
land improvements
10080 =25200*40%